OpenStudy (anonymous):

Prove the Power Rule using "n" factorization. I don't get how "n" factorization works. Prove the Power Rule using "n" factorization. I don't get how "n" factorization works. @Mathematics

OpenStudy (jamesj):

What is "n factorization" and what is the power rule?

OpenStudy (anonymous):

(a + b)^n The Power Rule is nx^n-1

OpenStudy (anonymous):

Ok to make it easier, I understand what factorials are. I know that 2! means 2*1 = 2.

OpenStudy (anonymous):

But I have no idea how to apply any of that factorial for "n" stuff to power rule.

OpenStudy (anonymous):

If you can explain, that will be awesome. This is the one thing that is throwing me off.

OpenStudy (anonymous):

I mis-read the question completely. Factorials would be nice for the expansion of (a+b)^n but I don't get what exactly needs to be proved.

OpenStudy (anonymous):

Ok wait one minute. Let me just copy and paste to you what my prof said. Don't worry, its brief.

OpenStudy (anonymous):

Quote Professor: Yes you do need it. I did not use the binomial theorem to prove the power rule. I used the factorization of A^n-B^n=(A-B)(A^{n-1}+A^{n-2}B+A^{n-3}B^2+.... +A B^{n-2}+B^{n-1}). You need to know this factorization.

OpenStudy (anonymous):

"it" is to know how to prove the Power rule.

OpenStudy (anonymous):

When I looked at that factorization, I was like "WTH..".

OpenStudy (anonymous):

Embarrassingly, I've read this proof recently but my mind has gone blank :(

OpenStudy (jamesj):

Yes. This identity is very useful and this is the one he's using: \[ a^n - b^n = (a - b)(a^{n-1} + a^{n-2}b + a^{n-3}b^2 + ... + a^2b^{n-3} + ab^{n-2} + b^{n-1}) \]

OpenStudy (anonymous):

Ok how the heck does this work. I am focused, and I'll try really hard to understand.

OpenStudy (jamesj):

Look at an example to make this clearer (hopefully!): For instance a^5 - b^5: \[ (a-b)(a^4 + a^3b + a^2b^2 + ab^3 + b^4) \] \[= a (a^4 + a^3b + a^2b^2 + ab^3 + b^4) \] \[ - b (a^4 + a^3b + a^2b^2 + ab^3 + b^4) \] \[ = a^5 + a^4b+a^3b^2 + a^2b^3 + ab^4 \] \[ - a^4b - a^3b^2 - a^2b^3 - ab^4 - b^5 \] Now see that nearly all the terms cancel out, leaving us with \[ = a^5 - b^5 \]

OpenStudy (anonymous):

One quick question: How did you get the long brackets?

OpenStudy (jamesj):

So hopefully that gives you comfort that this identity for general n works.

OpenStudy (anonymous):

One beside (a-b)

OpenStudy (anonymous):

The long bracket beside (a-b), how did you get that?

OpenStudy (jamesj):

I'm proving the identity, starting in this case with the right-hand side above and showing it is equal to the left-hand side

OpenStudy (anonymous):

Alright continue on please.

OpenStudy (anonymous):

Oh okay, so what you basically did was take (a-b)*(a-b)^4 to show (a-b)^5?

OpenStudy (jamesj):

No. I didn't do that at all.

OpenStudy (jamesj):

No induction use.

OpenStudy (jamesj):

Just explicit calculation

OpenStudy (jamesj):

Look at what I did for n = 5. Write out for yourself the identity for n = 6 so you feel comfortable with it.

OpenStudy (anonymous):

Then where did you get (a^4 + a^3b + a^2b^2 + ab^3 + b^4) from?

OpenStudy (anonymous):

This is what I was saying about the bracket beside (a-b)

OpenStudy (jamesj):

It is the term we obtain when we evaluate \[ \frac{a^5 - b^5}{a-b} \]

OpenStudy (jamesj):

So either you prove that it works by doing what I did above; or try the long division.

OpenStudy (anonymous):

Ok, so you are telling me to do: (a^6 - b^6)/(a-b)?

OpenStudy (jamesj):

Better: show that a^6 - b^6 = (a - b)(a^5 + a^4b + .... + b^5)

OpenStudy (precal):

Good job JamesJ. Sometimes it is hard for students to make the jump from a specific example to an abstract one. Students needs to realize proofs are abstract and apply to all cases.

OpenStudy (anonymous):

OHH Ok, I get it so how do I use this approach with Proving the Power Rule? I understand now that that is how you get that bracket^(whatever no.).

OpenStudy (jamesj):

Now then. To calculate the derivative of f(x) = x^n we need to evaluate the limit of the following difference ratio: \[ \frac{ f(x) - f(a)}{x-a} \] \[ = \frac{x^n - a^n}{x-a} \]

OpenStudy (anonymous):

Why not: f(x+h) - f(x) --- h?

OpenStudy (jamesj):

Same difference. It's actually slightly notationally easier with the notation I'm using, but they're otherwise identical. From my notation to yours: x --> x + h , a --> x

OpenStudy (jamesj):

but we'll use the form you're using. Write g = x + h. Then we want \[ \frac{ g^n - x^n}{g-x} \]

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

Now \[g^n - x^n = (g-x)(g^{n-1} + .... + x^{n-1}) \] hence the ratio is just \[ g^{n-1} + .... + x^{n-1} \]

OpenStudy (jamesj):

Notice that this sum has n terms where the power of g and the power of x sum to n-1.

OpenStudy (jamesj):

and all the coefficients of these monomials g^j x^(n-1-j) are 1.

OpenStudy (jamesj):

Now the limit of each term as h --> 0 is x^j . x^(n-1-j) = x^(n-1)

OpenStudy (jamesj):

therefore the sum of the n of them is \[ nx^{n-1} \]

OpenStudy (anonymous):

Why do you have those dots in b/ween "...."

OpenStudy (jamesj):

Because there are n of them in general, but I can't write down n terms in general explicitly without using some complicated notation. I.e., is it n = 2, or n = 17, or n = 256809?

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

But if you want to see it explicitly, here it is: \[ a^n - b^n \ = \ (a-b) \ \sum_{j=0}^{n-1} a^{n-1-j}b^j \]

OpenStudy (anonymous):

I get that. So you are doing n choose j,

OpenStudy (jamesj):

No

OpenStudy (precal):

Make the leap QRAwarrior from specific to abstract. JamesJ has done a great job of explaining it to you. :)

OpenStudy (jamesj):

There are no binomial coefficients

OpenStudy (jamesj):

So, formally, if \( f(x) = x^n \) then \[ f'(x) = \lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}{h} = \lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}{(x+h) - x} \]

OpenStudy (jamesj):

\[ = \lim_{h \rightarrow 0} \frac{ ((x+h) - x) \sum_{j=0}^{n-1} (x+h)^{n-1-j} x^j }{(x+h) - h} \]

OpenStudy (jamesj):

\[ = \lim_{ h \rightarrow 0} \ \ \sum_{j=0}^{n-1} (x+h)^{n-1-j} x^j \] \[ = \sum_{j=0}^{n-1} x^{n-1-j} x^j \] \[ = \sum_{j=0}^{n-1} x^{n-1} \] \[ = x^{n-1} \ \sum_{j=0}^{n-1} \ 1 \] \[ = n x^{n-1} \]